OpenStudy (loser66):

Please, help if a >b Prove \(a^2b

OpenStudy (anonymous):

Alrighty then.

OpenStudy (loser66):

@xapproachesinfinity help me, friend?

OpenStudy (loser66):

so far, I got a > b a^3 > b^3 --> a^3 -b^3 >0 --> \(\dfrac{a^3-b^3}{3}> 0\) and \(\dfrac{a^3-b^3}{3}+ab^2 -ab^2> 0\\\dfrac{a^3-b^3}{3}+ab^2>ab^2\)

OpenStudy (loser66):

but ab^2 is not > a^2b , so that I am stuck

OpenStudy (xapproachesinfinity):

hmm try to show that ab^2>a^2b

OpenStudy (mathmath333):

\(\large a>b\) \(\large b<a\) \(\large 0<a-b\) \(\large 0<(a-b)^2\) \(\large 0<(a^2-2ab+b^2)\) \(\large 0<(a^2-3ab+ab+b^2)\) \(\large 3ab<(a^2+ab+b^2)\) \(\large 3ab(a-b)<(a-b)(a^2+ab+b^2)\) \(\large 3ab(a-b)<(a^3-b^3)\) \(\large (a^2b-ab^2)<(\dfrac{a^3-b^3)}{3}\)

OpenStudy (loser66):

hahaha... how?? a > b a^2 > ab a^2b > ab^2 not conversely

OpenStudy (xapproachesinfinity):

eh he got the answer lol

OpenStudy (mathmath333):

cool

OpenStudy (loser66):

Yes, it 's perfect. Thank you

OpenStudy (mathmath333):

yw

OpenStudy (xapproachesinfinity):

I'm sure you go your way too

OpenStudy (xapproachesinfinity):

can*

OpenStudy (xapproachesinfinity):

what you said here is not true a>b then a^2>ab we have no idea what a is positive or negative

OpenStudy (loser66):

yes, you are right.

OpenStudy (xapproachesinfinity):

eh it is really hard to go your way! not knowing about a and b except that a>b

OpenStudy (loser66):

Hey, help me make the next problem clear, ok?

OpenStudy (xapproachesinfinity):

you got another problem hehe

OpenStudy (loser66):

I confused!!! the problem: Prove that in an ordered field , if \(\sqrt 2\) is a positive number whose square is 2, then \(\sqrt 2 <3/2\)

OpenStudy (loser66):

I don't see any mistake on my logic, but it leads me to nowhere

OpenStudy (loser66):

Suppose \(\sqrt 2 > 3/2\) then \(\sqrt 2 - 3/2 >0\) square a positive number gives us a positive number also. So that \((\sqrt 2 -3/2)^2>0\) it gives me \(2 -3\sqrt 2+9/4 >0\) that means \(17/4 > 3\sqrt 2\) or \(17/12 > \sqrt 2\) since both sides are >0, we square it , the sign doesn't change and I get 2.0006 > 2 :((((

OpenStudy (anonymous):

I'm really sleepy.

OpenStudy (loser66):

ok, go to bed, friend. :)

OpenStudy (xapproachesinfinity):

you started with \(\sqrt{2}>3/2 \) then you reached \(\sqrt{2}<17/12<3/2\) so contradiction here

OpenStudy (xapproachesinfinity):

the problem lies in that step!

OpenStudy (xapproachesinfinity):

you can stop in there and go back the conclusion that \(\sqrt{2}>3/2\) is not true \(\therefore \sqrt{2}<3/2\)

OpenStudy (loser66):

Gotttttttttttt you, haaahaa... I didn't apply brake on my run

OpenStudy (xapproachesinfinity):

yeah you were running fast down hill hehehe

OpenStudy (xapproachesinfinity):

number theory?

OpenStudy (loser66):

advance cal . prove everything :)

OpenStudy (xapproachesinfinity):

hehe! your course is dealing with pure proofs hehe

OpenStudy (xapproachesinfinity):

proves*

OpenStudy (mathmath333):

this is the general,lol \(\large \sqrt{2}<\dfrac{3}{2}\) \(\large \ 2 \times \sqrt{2}<3\) \(\large \sqrt{2 \times 4}<3\) \(\large \sqrt{8}<\sqrt{9}\) \(\large \ 8<9\)

OpenStudy (mathmath333):

cant understand that ordered field

OpenStudy (xapproachesinfinity):

hehe actually you next step is fine! since 2<2.0006 is not true therefore the assumption is not true

OpenStudy (loser66):

@xapproachesinfinity 2 < 2.0006 is true

OpenStudy (loser66):

@mathmath333 ordered field means <

OpenStudy (xapproachesinfinity):

eh my bad not paying the attention lol

OpenStudy (loser66):

It 's a field

OpenStudy (xapproachesinfinity):

exhausted hehe

OpenStudy (loser66):

so that it follows some rule.

OpenStudy (xapproachesinfinity):

you can't really go conversely! can you?

OpenStudy (loser66):

@xapproachesinfinity go to bed, friend. hahahaha

OpenStudy (mathmath333):

lol

OpenStudy (loser66):

I HAVE to use the rule of if a <b then a -b <0,

OpenStudy (loser66):

and manipulate everything base on it.

OpenStudy (loser66):

for example, if a^2 -b^2 <0, I must break it as (a+b)(a-b)<0 it <0 when one of them >0 and other <0 , so on and so on

OpenStudy (xapproachesinfinity):

eh i should get some rest haha!

OpenStudy (loser66):

sleeeeeeeep !! you are lucky. I have a bunch of homework due on Monday. I have to work out all of them.

OpenStudy (xapproachesinfinity):

i have a bunch too! but easy stuff you know hehe

OpenStudy (loser66):

:)

OpenStudy (anonymous):

We all should sleep. I'll bring the drugged popcorn.